HETAV PATEL

Grade 11,

In any triangle ABC if AB=2 ,BC=4,AC=3 AND D is the midpoint of BC then cos B=?,cosC=?

In any triangle ABC if AB=2 ,BC=4,AC=3 AND D is the midpoint of BC then cos B=?,cosC=?

Grade:11

1 Answers

Arun
6 years ago
cos B = (a2 + c2 –b2)/ 2 a*c
similarily cos C = (a2 + b2 –c2)/ 2 a*b
put the values, you wil get answers
where a = 4, b = 3, c = 2
hence cos B = 11/16
cos C = 21/24
hope it helps